A body vibrating with simple harmonic motion has a frequency of 4Hz and an amplitude of 0.15m. calculate
i) the maximum values of the acceleration and velocity
Ii) the acceleration and velocity at a point 0.09m from the equilibrium.
1
Expert's answer
2021-09-29T09:54:48-0400

Given:

"\omega=2\pi f=2*3.14*4\:\rm Hz=25\: rad/s"

"A=0.15\:\rm m"


The equation of motion of SHO:

"x=A\cos\omega t"

The velocity

"v=x'=-A\omega\sin\omega t\\=-A\omega\sqrt{1-\cos^2\omega t}=-\omega\sqrt{A^2-x^2}"

The acceleration

"a=v'=-A\omega^2\cos\omega t=-\omega^2x"

Hence,

i) the maximum values of the acceleration and velocity


"a_{\max}=\omega^2A=25^2*0.15=94.7\:\rm m/s^2"

"v_{\max}=\omega A=25*0.15=3.75\:\rm m/s"

ii) the acceleration and velocity at a point 0.09m from the equilibrium

"a=-\omega^2x=-25^2*0.09=-56.3\:\rm m/s^2"

"v=-25\sqrt{0.15^2-0.09^2}=-3.00\rm \: m/s"